OpenStudy (anonymous):

What is the size of the matrix that results in multiplying a 5 by 3 matrix by a 3 by 2 matrix? A. undefined B. 5 by 3 matrix C. 5 by 2 matrix D. 2 by 5 matrix

OpenStudy (jdoe0001):

\(\begin{bmatrix} \square &\square &\square \\ \square &\square &\square \\ \square &\square &\square \\ \square &\square &\square \\ \square &\square &\square \\ \end{bmatrix}\times \begin{bmatrix} \square &\square &\square \\ \square &\square &\square \\ \square &\square &\square \\ \end{bmatrix}\quad ?\)

OpenStudy (jdoe0001):

or you meant \(\begin{bmatrix} \square &\square &\square&\square &\square \\ \square &\square &\square&\square &\square \\ \square &\square &\square&\square &\square \\ \end{bmatrix}\times \begin{bmatrix} \square &\square &\square \\ \square &\square &\square \\ \square &\square &\square \\ \end{bmatrix}\quad ?\)